Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Box A contains 2 black chips. Box B contains 2 white chips. Box C contains 1 black

Expert replies
by BTGModeratorVI » Wed Feb 03, 2021 10:43 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Box A contains 2 black chips. Box B contains 2 white chips. Box C contains 1 black chip and 1 white chip. Ted chooses a box at random and then randomly selects a chip from that box. If the selected chip is black, what is the probability that the other chip in the same box is also black?

A) 1/5
B) 1/4
C) 1/3
D) 1/2
E) 2/3

Answer: E
Source: GMAT prep now
Join the discussion
Source: — Problem Solving |

BTGModeratorVI wrote: ↑
Wed Feb 03, 2021 10:43 am
Box A contains 2 black chips. Box B contains 2 white chips. Box C contains 1 black chip and 1 white chip. Ted chooses a box at random and then randomly selects a chip from that box. If the selected chip is black, what is the probability that the other chip in the same box is also black?

A) 1/5
B) 1/4
C) 1/3
D) 1/2
E) 2/3

Answer: E
Source: GMAT prep now
Let’s let:
B1 be one black chip in Box A.
B2 be the other black chip in Box A.
W1 be one white chip in Box B.
W2 be the other white chip in Box B.
B3 be the black chip in Box C
W3 be the white chip in Box C

There are 6 EQUALLY LIKELY outcomes:
Case i: Ted selects B1 from box A
Case ii: Ted selects B2 from box A
Case iii: Ted selects W1 from box B
Case iv: Ted selects W2 from box B
Case v: Ted selects B3 from box C
Case vi: Ted selects W3 from box C

If the selected chip is black, then we’re dealing with cases i, ii, or v
All 3 cases are EQUALLY LIKELY
In case i, the other chip is also black
In case ii, the other chip is also black
In case v, the other chip is white


So, of the 3 possible cases, 2 cases are such that the other chip in the box is black.

P(other chip is black) = 2/3

Answer: E

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

BTGModeratorVI wrote: ↑
Wed Feb 03, 2021 10:43 am
Box A contains 2 black chips. Box B contains 2 white chips. Box C contains 1 black chip and 1 white chip. Ted chooses a box at random and then randomly selects a chip from that box. If the selected chip is black, what is the probability that the other chip in the same box is also black?

A) 1/5
B) 1/4
C) 1/3
D) 1/2
E) 2/3

Answer: E
Solution:

Let’s look at the possible outcomes and their respective probabilities. Let’s let A, B, or C stand for the box that is selected, and let’s let X = the outcome that a black chip is drawn. We see that the probability of selecting a particular box is ⅓ and that the probability of drawing a black chip from box A is 1, from box B is 0, and from box C is ½.

For box A, the outcome (A, X) has a probability of 1/3 x 1 = 1/3.

For box B, the outcome (B, X) has a probability of 1/3 x 0 = 0 (i.e., we can’t draw a black chip).

For box C, the outcome (C, X) has a probability of 1/3 x ½ = 1/6.

Now, the only way that the other chip in the box is also black requires that we selected box A. Thus, we have the conditional probability statement “Given that a black chip was drawn, what is the probability that we chose box A?”.

The total probability that a black chip was drawn is 1/3 + 1/6 = 2/6 + 1/6 = 3/6 = 1/2. The probability that the box selected was box A will thus be 1/3 / 1/2 = 2/3.

Alternate Solution:

Since we know the selected chip is black, it is impossible for this chip to be drawn out of box B. Of the remaining two boxes, box A contains a greater number of black chips, therefore the probability that box A was selected should be greater than the probability that box C was selected. Since 2/3 is the only answer choice greater than 1/2, it must be the correct choice.

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion